OpenStudy (anonymous):

Find x and y so that the equation is true (x+y)^2 + (x+y)^2 = (x^2 + 2xy + y^2)

OpenStudy (anonymous):

it can be any number for x and y as long as the equation is true in the end

OpenStudy (anonymous):

are you sure that's the right equation? $$x^2+2xy+y^2=(x+y)^2$$so your question is equivalent to $$(x+y)^2+(x+y)^2=(x+y)^2\\\implies (x+y)^2=0\\\implies x=-y,y\in\mathbb{R}$$

OpenStudy (anonymous):

it is an equation I had to make. If you wait i can give you the list of equations

OpenStudy (anonymous):

Column A (x-y) (x+y) (y+x) (y-x) Column B (x^2 + 2xy + y^2) (x^2 - 2xy + y^2) (ax+b) (cy+d)

OpenStudy (anonymous):

there you go @oldrin.bataku

OpenStudy (anonymous):

if you could figure out a polynomial identity that works by combining 1 thing from column A and one from column B.

OpenStudy (anonymous):

x=0 and y=0

OpenStudy (anonymous):

besides 0 and 0 :/

OpenStudy (anonymous):

@raesal